### v1

@@ 3.3.2 Text Encoder
3.4 Cross-modal MLP
Based on the Text Encoder, we get the linguistic embeddings, including word and sentence embeddings. However, there is a significant semantic gap between linguistic embedding space and visual embedding space. Following MiniGPT-4 [55], we design a cross-modal MLP which consists of one hidden layer $L_{s}$ to effectively align the linguistic embedding space and visual embedding space, as shown in the green part in the bottom of Figure 1. Specifically, for each word embedding $f_{i}$ in $\mathcal{F}_{e}$, the sparse embedding can be obtained by adopting the cross-modal MLP $L_{s}$, i.e.,
@@ 3.4 Cross-modal MLP
3.5 Dense Attention
To enhance the visual features and provide additional cues for the mask decoder, we introduce the Dense Attention Module to fuse visual and linguistic features, as shown in the orange part in the top middle of Figure 1. Firstly, we concatenate the aligned sentence embeddings $f_{e^{\prime}}^{s}$ and sparse embeddings $\mathcal{F}_{sparse}=\{f_{j}\}_{j=1}^{L}$ together:
@@ 3.5 Dense Attention
In the next step, a similarity calculation between $\mathcal{F}_{fix}$ and visual embeddings $\mathcal{F}_{v}$ is performed to get spatial attention $\mathcal{F}_{sp}$. Then, the dot product is applied between $\mathcal{F}_{sp}$ and $\mathcal{F}_{fix}$ to get spatial language attention, i.e.,
@@
Then, we concatenate the spatial language attention $\mathcal{F}_{sl}$ and visual embeddings $\mathcal{F}_{v}$ and apply a convolutional layer to reduce the number of channels to match the channel dimension of visual embeddings, i.e.,
@@
3.6 Mask Decoder
The mask Decoder of the vanilla SAM leverages sparse embeddings (point and box) from the prompt encoder and dense embeddings (mask) from the SAM predictor to get final predictions. Following this principle, we construct sparse embeddings and dense embeddings with the same shape which encode useful visual and linguistic features from the Cross-Modal MLP and Dense Attention module. Then we put sparse embeddings $\mathcal{F}_{sparse}$ and dense embeddings $\mathcal{F}_{dense}$ together with visual embeddings $\mathcal{F}_{v}$ into the Mask Decoder to get the mask predictions, i.e.,
@@ 3.6 Mask Decoder
where $M$ is the output of the Mask Decoder. Finally, the Mask Decoder uses the output $M$ with the highest score as the predicted mask.
4 Experiments
4.1 Experiment Settings
4.1.1 Datasets
We conduct experiments on two challenging referring video object segmentation datasets: Ref-Youtube-VOS [31] and Ref-DAVIS17 [29]. Ref-Youtube-VOS is a large-scale referring video object segmentation dataset, which contains around 15K referring expressions for 3,900+ videos. Each video has pixel-level instance segmentation annotations for every fifth frame. Ref-DAVIS17 is built upon DAVIS17 [30] by providing the language description for each specific object in each video and contains 90 videos. We follow the common practice and use the default split for training and testing.
4.1.2 Evaluation Metrics
@@
The primary evaluation metrics of Ref-Youtube-VOS and Ref-DAVIS17 are the average of the region similarity ($\mathcal{J}$) and the contour accuracy ($\mathcal{F}$), denoted by $\mathcal{J}\&\mathcal{F}$. For Ref-Youtube-VOS, since the annotations of the validation set are not publicly available, we evaluate our method on the official challenge server [^1] https://codalab.lisn.upsaclay.fr/competitions/3282. As for Ref-DAVIS17, it is evaluated using the official evaluation code [^2] https://github.com/davisvideochallenge/davis2017-evaluation.
4.2 Implementation Details
4.2.1 Model Settings
@@
We use the image encoder of SAM, a ViT backbone [35] as our visual encoder. For the text encoder, we use the Hugging Face [56] implementation of T5-3b [53]. The Cross-Modal MLP consists of one hidden layer, which employs the rectified linear unit (ReLU) activation function. The convolution layer we use in Dense Attention applies a set of $512$ $1\times 1$ filters to the input feature maps, resulting in 256 output feature maps. We freeze the parameters of the Visual Encoder and Text Encoder during the entire training stage. We use the last stage features from the Visual Encoder as the input to Dense Attention Module and Mask Decoder. The dimension of the input embeddings (visual embeddings, sparse embeddings, and dense embeddings) to the Mask Decoder is $\mathcal{C}=256$.
4.2.2 Training and Inference Details
@@
Follow [4], our experiments follow the pre-training and fine-tuning process. we perform pre-training on the Ref-COCO dataset [57] for 15 epochs. In practice, we did not reach the optimal number of epochs, but it was sufficient for practical use, demonstrating the efficiency of our method. And then, we fine-tune it on Ref-Youtube-VOS for 4 epochs. The data augmentation includes random horizontal flip, random resize, random crop, and photometric distortion. We employ Dice loss and Focal loss as our primary loss functions. Our model is implemented in PyTorch. Training is conducted on 8 NVIDIA Tesla A100 GPUs.
During inference, the video frames are down-scaled to 360p. We directly output the predicted segmentation masks without using any post-processing tricks.
4.3 Main Results
Table 1: Results on Ref-DAVIS17.
@@ 4.3 Main Results
PerSAM [50]+Grounding Dino [25] 40.3 38.2 42.4
SAM-Track+Grounding Dino [21] 57.1 54.6 59.7
RefSAM 66.1 62.9 69.3
4.3.1 Ref-DAVIS17
@@
To have a closer look at the effectiveness of our method, we also compare our method with some recent SAM-based methods. Since there are no off-the-shelf SAM-based models for RVOS, existing SAM-based methods, such as SAM-Track [21] and PerSAM [50], can be adapted to RVOS task by providing bounding boxes in the first frame. Specifically, simple SAM-based RVOS baselines can be derived by combining SAM-Track or PerSAM with Grounding Dino [25], which is an object detector that can provide bounding boxes of objects in the first frame with referring expressions. As shown in Table 1, these naive solutions exhibit satisfactory performance, surpassing the CMSA and URVOS methods, yet significantly lagging behind our RefSAM model. Firstly, these methods employ a two-stage pipeline that can lead to a sub-optimal performance in the downstream segmentation task due to inaccurate bounding boxes detected by the object detector in the initial frame. Secondly, the absence of specific fine-tuning on the RVOS dataset poses a challenge, primarily due to the substantial model size and the inherent two-stage design. It is also noteworthy that the involvement of individual models for each sub-task also brings extra difficulties for model deployment. These findings underscore the efficacy of our RefSAM’s end-to-end design and parameter-efficient tuning strategy.
4.3.2 Ref-Youtube-VOS
Table 2: Results on Ref-Youtube-VOS.
@@ 4.3.2 Ref-Youtube-VOS
ReferFormer [4] 64.9 62.8 67.0
PerSAM [50]+Grounding Dino [25] 30.8 27.1 34.4
SAM-Track+Grounding Dino [21] 52.2 49.9 54.5
RefSAM 55.1 53.9 56.3
@@ 4.4 Ablation Study
4.4.1 Effect of the Key Modules
RefSAM employs a Cross-Modal MLP and Dense Attention for referring video object segmentation. Table 3 summarizes the results of our baseline model when excluding some of the key modules. In each experiment, we first pre-train the models on Ref-COCO for 15 epochs and then fine-tune them on Ref-Youtube-VOS for 4 epochs. We evaluate them on Ref-Davis. Compared to our baseline, without using the Cross-Modal MLP and Dense Attention will lead to a performance drop of 1.56 J&F-Mean and 2.38 J&F-Mean, respectively, validating their importance in aligning and fusing visual and language features.
Table 3: Ablation study of the key components of RefSAM.
@@ 4.4.1 Effect of the Key Modules
RefSAM 58.5 55.6 61.4
w/o Cross Modal MLP 56.9(-1.6) 55.3 60.6
w/o Dense Attention 56.1(-2.4) 53.4 58.8
4.4.2 Influence of Hyper-parameter Settings
@@
(a) Number of hidden layers in Cross-Modal MLP
(b) Linguistic feature in Dense Attention
(c) Model size of Visual Encoder
Figure 3: Ablation study of different module designs.
Figure 2 presents the results of different learning rate settings for the learnable modules of RefSAM, including the Cross-Modal MLP, the Dense Attention, and the Mask Decoder. As can be seen, RefSAM converges faster and better when the learning rate is set to 1e-4 for the Cross-Modal MLP and Dense Attention, and 1e-6 for the Mask Decoder.
Furthermore, we study the influence of the number of hidden layers in the Cross-Model MLP. As shown in Figure 2(a). As can be seen, using a single hidden layer delivers faster and better convergence, which is the default setting. In Figure 2(b), we investigate the choice of linguistic feature in Dense Attention. As can be seen, only using the word feature leads to a slightly better result, which is the default setting.
4.5 Influence of Model Size
We then investigate the influence of model size, specifically the performance of using different sizes of Visual Encoder, including ViT-B, ViT-L, and ViT-H. For each model, we pre-train it on Ref-COCO for 15 epochs, then fine-tune it on Ref-Youtube-VOS for 4 epochs, and evaluate it on Ref-DAVIS17. We do not use the data augmentation described in Sec. 4.2.2. The results are shown in Figure 2(c) and Table 4. The findings indicate that employing a larger backbone for visual feature extraction yields superior performance compared to a smaller one, attributed to the enhanced representation capacity of larger vision transformers. Furthermore, the results exhibit a consistent improvement in performance with increasing model size, highlighting the good scalability of our RefSAM model.
Table 4: Influence of the model size of Visual Encoder in RefSAM.

### v3

@@ III-C2 Text Encoder
III-D Cross-modal MLP
Based on the Text Encoder, we get the linguistic embeddings, including word and sentence embeddings. However, there is a significant semantic gap between linguistic embedding space and visual embedding space. Inspired by MiniGPT-4 [54], we design a cross-modal MLP $L_{s}$ which consists of one hidden layer to effectively align the linguistic embedding space and visual embedding space, as shown in the green part in the bottom of Figure 2. Specifically, for each word embedding $f_{i}$ in $\mathcal{F}_{e}$, the sparse embedding can be obtained by adopting the cross-modal MLP $L_{s}$, i.e.,
@@ III-D Cross-modal MLP
III-E Hierarchical Dense Attention
Figure 3: The structure of HDA. 1) The left part denotes the overall architecture of HDA. The inputs include the Intermediate Embeddings, Visual Embeddings, and Sparse Embeddings. The output is fine-grained Dense Embeddings. 2) The right part denotes the structure of Dense Attention for fusing hierarchical visual semantic information and sparse embeddings.
III-E1 Architecture
In our experiment, we discovered that the original pipeline of SAM, which solely utilizes the output of the Visual Encoder in the mask decoder, struggles to effectively identify tiny objects and leverage intricate text prompts. To address this isuee, we propose a novel HDA module, as shown in Figure 3, which leverages different levels of visual features and textual features to acquire fine-grain cross-modal embeddings. Specifically, we put intermediate embeddings $\mathcal{F}_{mid}=\{f^{mid}_{t}\}_{t=1}^{T}$, visual embeddings $\mathcal{F}_{v}=\{f_{t}\}_{t=1}^{T}$, and sparse embeddings $\mathcal{F}_{sparse}=\{f_{j}^{s}\}_{j=1}^{L}$ together with aligned sentence embeddings $f_{e^{\prime}}^{s}$ into the HDA module. For each feature map, we combine it with sparse embeddings and aligned embeddings in the Dense Attention (DA) module for pixel-level fusion. The output of each Dense Attention will be summed to generate the final Dense Embeddings. For intermediate feature maps, we use an extra $1\times 1$ convolutional layer to reduce its dimensionality before feeding them into Dense Attention.
III-E2 Dense Attention
To enhance the visual features and provide additional cues for the mask decoder, we propose the DA module for fusing visual and linguistic features at the pixel level, as shown in the right part of Figure 3. For example, given the visual embeddings $\mathcal{F}_{v}$, we firstly concatenate the aligned sentence embeddings $f_{e^{\prime}}^{s}$ and sparse embeddings $\mathcal{F}_{sparse}=\{f_{j}\}_{j=1}^{L}$ together:
@@ III-E2 Dense Attention
Next, we calculate the similarity between $\mathcal{F}_{fix}$ and $\mathcal{F}_{v}$ to get the attention map $\mathcal{F}_{sp}$. Then, we apply the dot product between $\mathcal{F}_{sp}$ and $\mathcal{F}_{fix}$ to get the attentive feature, i.e.,
@@
Then, we concatenate $\mathcal{F}_{sl}$ and $\mathcal{F}_{v}$ and apply a convolutional layer to reduce the number of channels to match the channel dimension of $\mathcal{F}_{v}$, i.e.,
@@
III-F Mask Decoder
The Mask Decoder of the vanilla SAM leverages sparse embeddings (point and box) from the prompt encoder and dense embeddings (mask) from the SAM predictor to get final predictions. Following this principle, we construct sparse embeddings and dense embeddings with the same shape which encode useful visual and linguistic features from the Cross-Modal MLP and HDA module. Then we put sparse embeddings $\mathcal{F}_{sparse}$ and dense embeddings $\mathcal{F}_{dense}$ together with visual embeddings $\mathcal{F}_{v}$ into the Mask Decoder to get the mask predictions, i.e.,
@@ III-F Mask Decoder
where $M$ is the output of the Mask Decoder. Finally, the Mask Decoder uses the output $M$ with the highest score as the final mask prediction.
III-G Implicit Tracking Module
In the context of the RVOS task, leveraging historical cues plays a crucial role. To this end, we devise an ITM $\mathcal{T}$ to transfer historical information to the subsequent frame. Specifically, in each frame, we extract the main mask token $E_{m}\in\mathbb{R}^{C_{v}}$, which is processed through the Mask Decoder to fuse with visual and text features, and then fed into $\mathcal{T}$ to obtain a track token $E_{track}$. $\mathcal{T}$ consists of two layers of feed-forward neural networks (FFNs) and includes a residual connection, a ReLU activation layer, and a final layer normalization. The shape of $E_{track}$ is same as $E_{m}$, i.e.,
$E_{track}=\mathcal{T}(E_{m})\in\mathbb{R}^{C_{v}},$ (8)
In current frame, We concatenate all mask tokens and the track token from the last frame in the Mask Decoder. Since $E_{m}$ encodes the mask of the target object in the current frame, the generated track token $E_{track}$ can provide valuable guidance for the location of the target object in the subsequent frame, i.e., $\mathcal{T}$ implicitly functions as a tracking module.
III-H Segmentation with Tracking
The training and inference processes are illustrated in Figure 1. For the first frame, the training and inference processes in the Mask Decoder follow the same approach as in Sec. III-F, and we get $E_{m}$ from the Mask Decoder, i.e.,
$M,E_{m}=Decoder(\mathcal{F}_{sparse},\mathcal{F}_{dense},\mathcal{F}_{v}),$ (9)
For the subsequent frame, we put sparse embeddings $\mathcal{F}_{sparse}$, dense embeddings $\mathcal{F}_{dense}$, and visual embeddings $\mathcal{F}_{v}$ together with the track token of the last frame $E_{track}$ into the Mask Decoder to get the mask prediction and $E_{m}$, i.e.,
$M,E_{m}=Decoder(\mathcal{F}_{sparse},\mathcal{F}_{dense},\mathcal{F}_{v},E_{track}).$ (10)
IV Experiments
IV-A Experiment Settings
IV-A1 Datasets
Our experimental evaluations were carried out on two demanding referring video object segmentation datasets: Refer-Youtube-VOS [27] and Ref-DAVIS17 [25]. Refer-Youtube-VOS is an extensive referring video object segmentation dataset that comprises approximately 15,000 referring expressions associated with more than 3,900 videos. Ref-DAVIS17 is an extension of the DAVIS17 dataset [26], where it enhances the dataset by providing language descriptions for each specific object present in the videos. It consists of a total of 90 videos. We follow the common practice and use the default split for training and testing.
IV-A2 Evaluation Metrics
@@
The main evaluation metrics used in Refer-Youtube-VOS and Ref-DAVIS17 are the average of region similarity ($\mathcal{J}$) and contour accuracy ($\mathcal{F}$), denoted as $\mathcal{J}\&\mathcal{F}$. Regarding Refer-Youtube-VOS, as the annotations for the validation set are not publicly accessible, we assess our method’s performance on the official challenge server  [^1] https://codalab.lisn.upsaclay.fr/competitions/3282. As for Ref-DAVIS17, it is evaluated using the official evaluation code [^2] https://github.com/davisvideochallenge/davis2017-evaluation.
IV-B Implementation Details
IV-B1 Model Details
@@
We use the image encoder of SAM, i.e., the ViT backbone [32], as our visual encoder. For the text encoder, we use the Hugging Face [55] implementation of T5-3b [50]. The Cross-Modal MLP consists of one hidden layer, which employs the rectified linear unit (ReLU) activation function. In HDA, we utilize the features from three intermediate layers of the visual encoder and the output feature maps of the visual encoder as the input. The convolution layer we use in DA applies a set of $512$ $1\times 1$ filters to the input feature maps, resulting in 256 output feature maps. ITM consists of two layers of feed-forward neural networks (FFNs) and includes a residual connection, a ReLU activation layer, and a final layer normalization. We freeze the parameters of the Text Encoder during the entire training stage. For the Visual Encoder, we tune the added adapter while freezing the other parameters throughout the entire training stage. We add the adapter to the latter half of the transformer blocks in SAM’s image encoder. We use the output feature maps from the Visual Encoder as the input to Dense Attention Module and Mask Decoder. The dimension of the input embeddings (including visual embeddings, sparse embeddings, and dense embeddings) to the Mask Decoder is $\mathcal{C}=256$.
IV-B2 Training and Inference Details
@@
Following [4, 56], our experiments contain the pre-training and fine-tuning process. As in SOC[56], we pre-train on the RefCOCO dataset for 40 epochs. Then, we fine-tune it on Refer-Youtube-VOS for 4 epochs, and we randomly sample $N=3$ frames in each batch size for training the ITM module. The data augmentation includes random horizontal flip, random resize, random crop, and photometric distortion. The learning rate is set to 1e-4 for the Cross-Modal MLP, HDA, and DA, 1e-6 for the Mask Decoder, and 1e-5 for the Adapter. We employ Dice loss and Focal loss as our primary loss functions. Our model is implemented in PyTorch and trained on 8 NVIDIA Tesla A100 GPUs. During inference, the video frames are down-scaled to 360p.
IV-C Main Results
TABLE I: Results on Ref-DAVIS17.
@@ IV-C Main Results
SgMg[58] 63.3 60.6 66.0
SOC[56] 64.2 61.0 67.4
FTVR[59] 64.2 61.1 67.2
OnlineRefer[60] 64.8 61.6 67.7
PerSAM[45]+Grounding Dino 40.3 38.2 42.4
SAM-Track[17]+Grounding Dino 57.1 54.6 59.7
RefSAM 71.9 68.4 75.5
IV-C1 Ref-DAVIS17
@@
To have a closer look at the effectiveness of our method, we also compare our method with some recent SAM-based methods. Since there are no off-the-shelf SAM-based models for RVOS, existing SAM-based methods, such as SAM-Track [17] and PerSAM [45], can be adapted to RVOS task by providing bounding boxes in the first frame. Specifically, simple SAM-based RVOS baselines can be derived by combining SAM-Track or PerSAM with Grounding Dino [21], which is an object detector that can provide bounding boxes of objects in the first frame with referring expressions. As shown in Table I, these naive solutions exhibit satisfactory performance, surpassing the CMSA and URVOS methods, yet significantly lagging behind our RefSAM model. Firstly, these methods employ a two-stage pipeline that can lead to a sub-optimal performance in the downstream segmentation task due to inaccurate bounding boxes detected by the object detector in the initial frame. Secondly, the absence of specific fine-tuning on the RVOS dataset poses a challenge, primarily due to the substantial model size and the inherent two-stage design. It is also noteworthy that the involvement of individual models for each sub-task also brings extra difficulties for model deployment. These findings underscore the efficacy of our RefSAM’s end-to-end design and parameter-efficient tuning strategy.
Finally, to demonstrate that fine-tuning SAM is superior to directly using SAM for post-processing, we apply SAM to the post-processing of FTVR’s inference results. For each inference result from FTVR, we extract the result’s bounding box and input it into SAM for more fine-grained segmentation. Using SAM for post-processing the results of FTVR improves the J&F score by 2.5 points, raising it to 66.7. However, this score is still 5.2 points lower than RefSAM, demonstrating the efficacy of fine-tuning SAM.
IV-C2 Refer-Youtube-VOS
TABLE II: Results on Refer-Youtube-VOS.
@@ IV-C2 Refer-Youtube-VOS
SgMg [58] 65.7 63.9 67.4
SOC [56] 66.0 64.1 67.9
PerSAM [45]+Grounding Dino 30.8 27.1 34.4
SAM-Track[17]+Grounding Dino 52.2 49.9 54.5
RefSAM 67.6 65.8 69.4
@@ IV-D Ablation Study
IV-D1 Effect of the Key Modules
TABLE III: Ablation study of the key components of RefSAM.
Cross-modal MLP DA HDA Refer-Youtube-VOS Ref-DAVIS17
@@ IV-D1 Effect of the Key Modules
48.5 47.8 49.1 57.1 53.9 60.3
49.7 49.0 50.3 57.8 54.8 60.9
51.8 51.9 51.8 59.1 56.2 61.9
50.2 49.8 50.6 59.3 56.6 62.1
52.2 52.0 52.5 60.3 57.3 63.3
RefSAM employs Cross-Modal MLP, HDA, and ITM for referring video object segmentation. Since ITM requires sampling $N$ frames for each batch size, which differs from the other modules in the experiment, we first investigate the effects of other key components by isolating ITM. Table III summarizes the results on Refer-Youtube-VOS and Ref-DAVIS17. DA means we only use the visual embeddings for Dense Attention. HDA means we use both visual embeddings and intermediate embeddings for Hierarchical Dense Attention. We use the data augmentation described in Sec. IV-B2. As shown in the example of Ref-DAVIS17, using both DA and Cross-modal MLP instead of HDA and Cross-modal MLP results in a performance drop of 1.0 in the J&F-Mean score. Additionally, using only the Cross-modal MLP or DA results in performance drops of 3.2 and 2.5 in the J&F-Mean score, respectively, compared to the best choice. On the other hand, using only HDA increases the performance by 1.3 compared to using only DA, but it still underperforms compared to using both HDA and Cross-modal MLP. On Refer-Youtube-VOS, using only HDA outperforms both Cross-modal MLP and DA. These results validate the importance of Cross-Modal MLP and HDA in aligning and fusing visual and language features.
TABLE IV: Ablation study of ITM.
[…]
Sing frame (Tab. III) 60.3 57.3 63.3
Three frames 59.6 56.5 62.7
IoU token 61.6 58.8 64.4
All mask tokens 61.9 58.9 64.8
Main mask token 62.2 58.9 65.5
Main mask token 62.5 59.5 65.4
Next, we ablate the design choice of ITM. The mask decoder of SAM outputs an IoU token, a main mask token, and three-scale mask tokens. We chose the IoU token, main mask token, and all four mask tokens separately as the input of ITM. For comparison, we conducted experiments with two control groups without ITM: one using a single frame and another using three frames as input. For all experiments with ITM, we utilize three frames in each batch size, employ Cross-modal MLP and HDA, and use the data augmentation described in Sec. IV-B2. The results on Ref-DAVIS17 are shown in Table IV. As can be seen, relying solely on three frames without incorporating other tracking or memory modules can decrease performance. This may be due to the model learning specialized knowledge from different frames instead of general knowledge, thus reducing the model’s generalization ability. Then, utilizing the main mask token as the input of ITM achieves the best performance with a J&F-Mean of 62.2, surpassing both the IoU token and all mask tokens. It is reasonable since the main mask token has a better capability of understanding general objects and can provide more accurate location information for a complete object than the other token that either emphasizes different granularities of objects or is responsible for predicting the IoU score. Furthermore, we investigate the structure design of ITM, i.e., with or without using the residual connection. The results in the last two rows show that utilizing a residual connection in ITM leads to an increase of 0.3 J&F-Mean.
IV-D2 Influence of Hyper-parameters
@@
Figure 4 presents the results of different learning rate settings for the learnable modules of RefSAM, including the Cross-Modal MLP, DA, and the Mask Decoder. As can be seen, RefSAM converges faster and better when the learning rate is set to 1e-4 for the Cross-Modal MLP and DA, and 1e-6 for the Mask Decoder. Furthermore, we study the influence of the number of hidden layers in the Cross-Model MLP. As shown in Figure 5, using a single hidden layer delivers faster and better convergence, which is the default setting.
Figure 5: The influence of the number of hidden layers in Cross-Modal MLP.
IV-D3 Impact of Different Text Encoders
We also study the impact of different text encoders in RefSAM, including 1) T5-3b [50], Roberta-base [61], and Clip-large [62]. All experiments use the data augmentation described in Sec. IV-B2 and the model employs Cross-modal MLP and HDA. The results on Ref-DAVIS17 are shown in V. As can be seen, RefSAM with the T5-3b text encoder performs better than the others, yielding an improvement of 1.6 J&F-Mean compared to Roberta-base and a more significant increase of 2.0 J&F-Mean over Clip-large.
TABLE V: Ablation study of the Text Encoder.